5. DEF Company makes bowling balls. Beginning finished goods inventory consisted of 16,000 balls with a total value of $96,000. During the month, there were 64,000 balls manufactured at a cost of $5 per balls and 70,000 balls were sold. Using the average costing method, determine the value of ending inventory?
Answer #1

Average cost per unit = (96000+64000*5)/(16000+64000)

= $5.2 per unit

Value of ending inventory = 10000*5.2 = $52000